OpenStudy (anonymous):

The perimeter of a rectangle is 96 ft. The ratio of its length to its width is 7 : 5. What are the dimensions of the rectangle?

OpenStudy (nurali):

Length = 7x Width = 5x Perimeter = 2*length + 2*width 96ft = 2*7x + 2*5x 96 ft = 14x + 10x 96ft = 24x 24x = 96ft x = 4ft Length = 4*7 = 28ft Width = 4*5 = 20ft. Hope this helps.

OpenStudy (anonymous):

Thanks

OpenStudy (nurali):

My Pleasure.